OpenStudy (anonymous):

integral (sinƟ + sinƟ tan^2 Ɵ)/sec^2 Ɵ, dƟ from 0 to pi/3

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/3} \frac{ \sin \Theta +\sin \Theta \tan^2 \Theta }{ \sec^2 \Theta } d \Theta \]

OpenStudy (anonymous):

please solve for me

OpenStudy (loser66):

factor sin theta out you have 1+ tan^2 inside which is sec^2 and cancel out with the denominator, so far you just have sin theta, how is it hard?

OpenStudy (anonymous):

hmmmm what about the 0 to pi/3 how does that play in

OpenStudy (anonymous):

sweet thanks :P